Shift report 20011225 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Dec 25 2001 - 23:54:32 EST

  • Next message: Hironori Ito: "Re: value of shift???"

    ------------------------------------------------------------
    Date: Tue Dec 25 23:54:32 2001, Shift: 16:00-24:00
    Supervisor: hiro, with: 
    Log book#11 pages 91 - 91
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam for the most of the shift (Finally!!!)
    2.  Data Run: 6148, 6149, 6150
    3.  Communication to tsir is dead.  
    I tried tuning off and on 1458B and tsir. But,
    I still can not communicate with 1458B.  It needs access to
    IR probably.  Therefore, for the above runs, the back platform is off since HVs of T3, T4, T5 and Rich are off.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Dec 25 2001 - 23:55:16 EST